OpenStudy (anonymous):

find the value of each variable 3x-2y+z=1 -x+y-z=2 5x+2y+10z=39

OpenStudy (campbell_st):

eliminate z first Equation 1 plus equation 2 2x - y = 3 (4) Equation 2 x 10 + equation 3 -5x + 12y = 59 (5) equation(4) x 12 + equation (5) to eliminate y 24x - 12y = 36 -5x + 12y = 59 19x = 95 x = 5 substitute into equation (4) to find y 2(5) - y = 3 y = 7 substitute into equation 2 to find z - 5 + 7 - z = 2 2 - z = 2 z = 0 the solution is x= 5, y = 7 and z = 0

OpenStudy (anonymous):

i had z=0 x=5 and y=-2 but now i get it

OpenStudy (campbell_st):

look at the 1st equation 3x - 2y + z = 1 15 + 4 + 0 = 1 not quite

OpenStudy (campbell_st):

where as 15 - 14 + 0 = 1 true

OpenStudy (anonymous):

yeah i realized that i understand the way its set up but i get confused with the substitution but thank you!!

OpenStudy (campbell_st):

hope it helps